# Accelerating Residual Reinforcement Learning with Uncertainty Estimation

### v1

@@ V-E Image Based Experiments
The results of our algorithm for the image-based Can Task is shown in Fig 6. Given that distance-to-data is a less reliable metric in high dimensional input spaces such as images, we instead used ensemble variance to quantify uncertainty. We explored two strategies: first, using ensemble variance to enhance the performance of a single policy, and second, leveraging it to improve an ensemble of policies. Our method was compared against DPPO, the strongest baseline for finetuning diffusion policies, and Residual Policy Learning (RPL) without uncertainty estimation. The results indicate that the ensemble-based approach starts with strong performance and avoids early dips, likely due to the robustness offered by ensembling. By contrast, the naive RPL strategy initially crashes to 0% performance after the initial online learning steps, though it recovers over time. DPPO, as observed previously, maintains steady but slow improvement. These findings show our approach also applies to image-based environments.
Figure 7: We present a novel approach for accelerating Residual RL. The learned policies were deployed in real world for the pick and place can task where the robot must pick the can from one side of the table and place it in the blue bin
Figure 8: Results for the policy deployment in real world. RL policies are robust in sim2real transfer even without any domain randomization.
@@ V-F Real World Experiments
Additionally, we deployed our learned policy for the Can task in the real world in a zero-shot setup as shown in Fig 7. The task is setup similarly to the simulated version, where the robot has to pick up the can from one side of the table and place it in the blue bin on the other side of the table. We used Grounded SAM [29] and Foundation Pose [30] to get the real-time state of the can object. We evaluated four policies in our real-world setup: 1) GMM base, 2) GMM base + Residual, 3) Diffusion base, and 4) Diffusion base + Residual. We compare the performance in terms of task progression divided into three stages: contacting, grasping, and placing. We note the performance of each policy on 10 trials in Fig 8. The performance of policies learned with Residual RL is able to nearly retain their original performance in simulation without any domain randomization, whereas the base policies struggled. This supports the observation that policies trained using RL tend to be more robust than behavior cloning-based policies, as they benefit from richer interaction with the environment.
@@ V-G Ablations
We ablated our proposed approach with different decaying strategies for the uncertainty threshold in Sec. V-G1, with different decay rates in Sec. V-G2, and for other kitchen environments in Sec. V-G3. All ablations are performed with the Diffusion base policy for the Can task.
@@ V-G1 Threshold Decay Strategy
Figure 10: Ablations for different decay rates.
Figure 11: Ablations for the other kitchen environments that contain undirected demonstrations.
@@ V-G2 Threshold Decay Rate
We evaluated our algorithm with different decay rates in Figure 11. Lower rates resulted in aggressive exploration, causing performance dips that were hard to recover from, while higher rates slowed the convergence. We need to balance decay rates for effective exploration without hindering convergence. However, our approach is not too sensitive to the chosen decay rate, as the performance is similar for decay rates ranging between 300k and 500k.
@@ V-G3 Kitchen Environments
We tested our method on the other two variations of the kitchen environments, Kitchen Partial and Kitchen Mixed. As seen in Figure 11, our method is not able to outperform the base policy. One key assumption we make in our algorithm is that when the base policy is certain, it will also be correct. This assumption does not hold true for these two environments: the base policies for these two environments are trained on the random play data, so the confidence and correctness of the policy are not strongly correlated, resulting in poorer performance.
@@ VI Conclusion and Future Work
We propose two improvements to the Residual RL framework to accelerate the learning with stochastic base policies. First, we use uncertainty estimates of the base policy to constrain the exploration of residual learning. Second, we adapt Residual RL to handle stochastic base policies by proposing an asymmetric actor-critic approach in which the critic observes the combined action, while the actor predicts only the residual action. Our method outperforms various baselines on tasks from Robosuite and the D4RL Benchmark. While our proposed method demonstrates strong performance, it would also benefit from a more robust epistemic uncertainty metric. We believe that, with reliable uncertainty metrics, our approach could also be applied to larger models including robot foundation models.
@@ Acknowledgments
This work was supported by NASA-Kennedy under grant GR5227163. SV was supported by the Office of Naval Research (ONR) under REPRISM MURI N000142412603 and ONR grant #N00014-22-1-2592, as well as by the National Science Foundation (NSF) via grant #1955361. Partial funding for SV was also provided by the Robotics and AI Institute. We would also like to thank Ondrej Biza for helpful discussions around the baselines and Ivy He for her help with real robot experiments.
APPENDIX
VI-A Environments
The observation space and the action space of each environment used are below -
@@ VI-A Environments
    Lift - 19 dim observation space with object state and robot end effector state. Action space is 7 DoF end effector pose.
@@
    Can - 23 dim observation space with object state and robot end effector state. Action space is 7 DoF end effector pose.
@@
    Square - 23 dim observation space with object state and robot end effector state. Action space is 7 DoF end effector pose.
@@
    Kitchen - 60 dim observation space with all object states and velocities, and robot joint states and angular velocity. The action space is 9 DoF joint angular velocity and gripper linear velocity.
VI-B Base Policies
VI-B1 GMM Base Policy
We use Robomimic [8] to train Gaussian Mixture Model based policies with a Recurrent Neural Network backbone.
VI-B2 Diffusion policy
We used the same base policies from DPPO [12] to keep the comparisons consistent. The Diffusion policies are trained with an action horizon of 1 and 20 denoising steps.
VI-C Hyperparameters
We used the same hyperparameters in each environment for both GMM-based and Diffusion-based. We keep the same parameters for actor and critic for Robosuite environments and use the advised hyperparameters from the DPPO paper for the kitchen environment. Hyperparameter details can be found in Table I. The uncertainty threshold value $U$ and decay rate values for our proposed approach with distance-to-data metric can be found in Table II, and with ensemble variance can be found in Table III.
@@ VI-C Hyperparameters
TABLE II: Uncertainty threshold $U$ and Decay Rate values for distance-to-data method
Environment Base Policy U Decay Rate
@@
TABLE III: Uncertainty threshold $U$ and Decay Rate values for ensemble variance
VI-D Tuning Policy Decorator
Policy Decorator  [14] has two hyperparameters namely, $\alpha$ the residual bound which scales the residual action to limit exploration and $H$ to schedule the exploration progressively. According to their paper, the $\alpha$ value is set close to the action scale of demonstration data while it is advised to keep $H$ large as a safe choice. We present the hyperparameter values we used in our sweep in Table IV. The authors perform an ablation with DPPO for the square task in their appendix, and we received the hyperparameters from the authors for the same. After looking at their implementation, we observed that they train their RL policies with an expanded action space with an action horizon while we implement the RL policies in the original form with a single action.
@@ VI-D Tuning Policy Decorator
TABLE IV: Residual bound $\alpha$ and Progressive Exploration schedule $H$ for Policy Decorator

### v2

@@ V-E Image Based Experiments
We evaluated our algorithm on the Can task with image observations. Given that distance-to-data is a less reliable metric in high dimensional input spaces such as images, we instead used ensemble variance to quantify uncertainty. We explored two strategies: first, using ensemble variance to enhance the performance of a single policy, and second, leveraging it to improve an ensemble of policies. Our method was compared against DPPO, the strongest baseline for finetuning diffusion policies, and Residual Policy Learning (RPL) without uncertainty estimation. The results of our algorithm on the image-based Can Task are shown in Fig 6. The results indicate that the ensemble-based approach starts with strong performance and avoids early dips, likely due to the robustness offered by ensembling. In contrast, the naive RPL strategy initially crashes to 0% performance after the initial online learning steps, though it recovers over time. DPPO, as observed previously, maintains steady but slow improvement. These findings show our approach also performs well in image-based environments.
Figure 7: Sim-to-Real setup for the Robomimic Can task.
Figure 8: Results for the policy deployment in real world.
@@ V-F Real World Experiments
We deployed our learned policy for the Can task in the real world in a zero-shot setup as shown in Fig 7. The task is set up similarly to the simulated version, where the robot has to pick up the can from one side of the table and place it in the blue bin on the other side of the table. We used Grounded SAM [22] and Foundation Pose [28] to get the real-time state of the can object. We evaluated four policies in our real-world setup: 1) GMM base, 2) GMM base + Residual, 3) Diffusion base, and 4) Diffusion base + Residual. We compare the performance in terms of task progression divided into three stages: contacting, grasping, and placing. Fig 8 shows the performance of each policy across 10 trials. Policies learned with Residual RL retain nearly all of their original performance in simulation without any domain randomization, whereas base policies struggled. This supports the observation that policies trained using RL tend to be more robust than behavior cloning policies, as they benefit from richer interaction with the environment.
@@ V-G1 Threshold Decay Strategy
Figure 10: Higher decay rates result in slower convergence whereas lower decay rates results in aggressive exploration.
Figure 11: Ensemble variance is a better uncertainty metric for Kitchen Partial and Kitchen Mixed environments as the training data also contains random trajectories.
@@ V-G2 Threshold Decay Rate
We evaluated our algorithm with different decay rates in Figure 11. Lower rates resulted in aggressive exploration, causing performance dips that were hard to recover from, while higher rates slowed the convergence. We need to balance decay rates for effective exploration without hindering convergence. However, our approach is robust to the chosen decay rate, as the performance is similar for decay rates ranging between 300k and 500k.
@@ V-G3 Kitchen Environments
We tested our method on the other two variations of the kitchen environments, Kitchen Partial and Kitchen Mixed. As seen in Figure 11, our method with the distance to data metric is not able to outperform the base policy. One key assumption we make in our algorithm is that when the base policy is certain, it will also be correct. This assumption does not hold true for these two environments: the base policies for these two environments are trained on the random play data, so the confidence and correctness of the policy are not strongly correlated, resulting in poorer performance. However, ensemble variance is able to provide a better uncertainty metric when random play data is involved and results in improved performance.
@@ VI Conclusion and Future Work
We propose two improvements to the Residual RL framework to accelerate the learning with stochastic base policies. First, we use uncertainty estimates of the base policy to constrain exploration in residual learning. Second, we adapt Residual RL to handle stochastic base policies by proposing an asymmetric actor-critic approach in which the critic observes the combined action, while the actor predicts only the residual action. While our proposed method demonstrates strong performance, it would also benefit from a more robust epistemic uncertainty metric. We believe that, with reliable uncertainty metrics, our approach could also be applied to larger models including robot foundation models. In the future, we would also like to dynamically set the values of uncertainty threshold decay based on the performance of the base policy.
@@ Acknowledgments
LD was supported by the NASA-Kennedy award “Efficient and High Quality Space Robotic VR Teleoperation through Neural Scene and Object Reconstruction.” under award number 80NSSC23M0075. SV and GK was supported by the Office of Naval Research (ONR) under REPRISM MURI N000142412603 and ONR grant #N00014-22-1-2592, as well as by the National Science Foundation (NSF) via grant #1955361. Partial funding for SV and GK was also provided by the Robotics and AI Institute. We would also like to thank Ondrej Biza and Ivy He for helpful discussions.
A. Environments
The observation space of each environment used is below:
@@ A. Environments
    Lift - 19 dim obs space with object state and robot eef.
@@
    Can - 23 dim obs space with object state and robot eef.
@@
    Square - 23 dim obs space with object state and robot eef.
@@
    Kitchen - 60 dim obs space with all object states and velocities, and robot joint states and angular velocity.
Action space for all Robosuite environments is 7 DoF end effector pose while for Kitchen environment is the 9 DoF joint angular velocity and gripper linear velocity.
B. Base Policies
GMM Base Policy - We use Robomimic [17] to train Gaussian Mixture Model based policies with a Recurrent Neural Network backbone.
Diffusion policy - We used the same base policies from DPPO [21] to keep the comparisons consistent. The Diffusion policies are trained with an action horizon of 1 and 20 denoising steps.
C. Hyperparameters
We used the same hyperparameters in each environment for both GMM-based and Diffusion-based. We keep the same parameters for actor and critic for Robosuite environments and use the advised hyperparameters from the DPPO paper for the kitchen environment. Hyperparameter details can be found in Table I. The uncertainty threshold value $U$ and decay rate values for our proposed approach with distance-to-data and ensemble variance metric can be found in Table II.
@@ C. Hyperparameters
Ensemble variance
@@
TABLE II: Uncertainty threshold $U$ and Decay Rate values for distance-to-data and ensemble variance.
D. Tuning Policy Decorator
Policy Decorator  [29] has two hyperparameters namely, $\alpha$ the residual bound which scales the residual action to limit exploration and $H$ to schedule the exploration progressively. According to their paper, the $\alpha$ value is set close to the action scale of demonstration data while it is advised to keep $H$ large as a safe choice. We present the hyperparameter values we used in our sweep in Table III. The authors perform an ablation with DPPO for the square task in their appendix, and we received the hyperparameters from the authors for the same. After looking at their implementation, we observed that they train their RL policies with an expanded action space with an action horizon while we implement the RL policies in the original form with a single action.
@@ D. Tuning Policy Decorator
TABLE III: Residual bound $\alpha$ and Progressive Exploration schedule $H$ for Policy Decorator.